# Towards An Analytical Framework for Dynamic Potential Games

### v2

@@ 1 Introduction
Throughout the paper, unless otherwise stated, proofs of theorems are deferred to Section 5. The proofs of certain technical lemmas can be found in Appendix A.
@@ 3 Distributed Game as Markov Potential Game
This crowd motion game has been studied in [4] for pedestrian motion modelling, in [10] for animal flocking, and [9] for synchronization in neuroscience. We emphasize that the form of game here generalizes those in [4, 10, 9]: condition (3.5) removes the homogeneous and the symmetric interaction conditions imposed in their analysis. In particular, it allows each player to have different individual preferences specified by $c_{i}$ and $\ell_{i}$, as well as to be differently influenced by other players’ positions as indicated by the kernels $h_{ij}$ and $\ell_{ij}$.
@@ 4 Characterization of Differentiable Potential Game
We then present two lemmas regarding the linear derivative of $f:\pi^{(N)}\rightarrow{\mathbb{R}}$, which are crucial for proving Theorem 4.3. Lemma 4.1 shows that if $f$ has a linear derivative in $\pi_{i}$, then $f$ is differentiable along any line segment within $\pi_{i}$. Additionally, Lemma 4.2 shows that if $f$ is differentiable concerning all unilateral deviations, then $f$ is differentiable with respect to perturbations of all players’ policies, and the derivative can be computed via a chain rule.
@@
We now show that if the value functions of a game are sufficiently regular and have a symmetric Jacobian, then this game is a potential game, and its potential function can be constructed via the linear derivative of its value function.
@@ 4.1 Probabilistic Characterization of Continuous-time Potential Game
where $B^{\phi}(t,x)\coloneqq B(t,x,\phi(t,x))$, $(\partial_{a_{i}}B^{\phi})[\phi^{\prime}_{i}](t,x)\coloneqq(\partial_{a_{i}}B)(t,x,\phi(t,x))\phi^{\prime}_{i}(t,x)$, and $\Sigma^{\phi}(t,x)$ and $(\partial_{a_{i}}\Sigma^{\phi})[\phi^{\prime}_{i}](t,x)$ are defined similarly. Here the differentiations are taken componentwise, i.e., $((\partial_{x}B^{\phi})(s,X^{t,x,\phi}_{s})Y_{s})_{\ell}=\sum_{j=1}^{n_{x}}\partial_{x_{j}}B^{\phi}_{\ell}(s,X^{t,x,\phi}_{s})(Y_{s})_{j}$ for all $\ell=1,\ldots,n_{x}$. Equation (4.6) is the sensitivity equation of (4.5) with respect to $\phi_{i}$ (see Lemma 5.1).
@@
and $F_{\Sigma}(\cdot,y_{1},y_{2},\phi^{\prime}_{i},\phi^{\prime\prime}_{j})$ is defined similarly. Equation (4.7) is the sensitivity equation of (4.5) with respect to $\phi_{i}$ and $\phi_{j}$ (see Lemma 5.2).
@@ 4.2 PDE Characterization of Continuous-time Potential Games
4.3 Linear Quadratic Differentiable Potential Game
@@
The following theorem proves the equivalence of Theorems 4.4 and 4.5 for LQ games.
@@ 5 Proofs of Main Results
The proof consists of first showing that the solutions to (4.6) and (4.7) characterize the derivatives of $X^{t,x,\phi}$ with respect to policies (see Lemmas 5.1 and 5.2), and then proving (4.9) and (4.10) characterize the derivatives of $\alpha^{t,x,\phi}$ with respect to policies (see Lemma 5.3). The proofs of Lemmas 5.1, 5.2 and 5.3 are given in Appendix A.2.
Lemma 5.1.
  1. (1)
  2. (2)
Lemma 5.2.
  1. (1)
  2. (2)
  3. (3)
Lemma 5.3.
  1. (1)
  2. (2)
Proofs of Theorem 4.4 Items 1 and 2.
Fix $\phi\in\pi^{(N)}$. The linearity of $\frac{\delta V^{t,x}_{i}}{\delta\phi_{j}}(\phi;\cdot):\operatorname*{span}(\pi_{j})\rightarrow{\mathbb{R}}$ follows from the linearity of $\frac{\delta X^{t,x}}{\delta\phi_{j}}(\phi;\cdot)$ and $\frac{\delta\alpha^{t,x}}{\delta\phi_{j}}(\phi;\cdot)$ shown in Lemmas 5.1 and 5.3. Let $\phi^{\prime}_{j}\in\pi_{j}$ and for each $\varepsilon\in[0,1)$, let $\phi^{\varepsilon}=(\phi_{j}+\varepsilon(\phi^{\prime}_{j}-\phi_{j}),\phi_{-j})$. Recall that
Then [25, Theorem 9, p. 97], the $\mathcal{S}^{\infty}$-differentiability of $X^{t,x,\phi^{\varepsilon}}$ and the $\mathcal{H}^{\infty}$-differentiability of $\alpha^{t,x,\phi^{\varepsilon}}$ (see Lemmas 5.1, 5.2 and 5.3) imply that $\frac{{\mathrm{d}}}{{\mathrm{d}}\varepsilon}V^{t,x}_{i}(\phi^{\varepsilon})|_{\varepsilon=0}=\frac{\delta V^{t,x}_{i}}{\delta\phi_{j}}(\phi;\phi^{\prime}_{j}-\phi_{j})$, hence Theorem 4.4 Item 1.
Proof of Theorem 4.4 Item 3.
$\|X^{t,x,\phi^{r,\varepsilon}}\|_{\mathcal{S}^{q}},\left\|\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi^{r,\varepsilon};\phi^{\prime}_{i})\right\|_{\mathcal{S}^{q}},\left\|\frac{\delta X^{t,x}}{\delta\phi_{j}}(\phi^{r,\varepsilon};\phi^{\prime\prime}_{j})\right\|_{\mathcal{S}^{q}},\left\|\frac{\delta^{2}X^{t,x}}{\delta\phi_{i}\phi_{j}}(\phi^{r,\varepsilon};\phi^{\prime}_{i},\phi^{\prime\prime}_{j})\right\|_{\mathcal{S}^{q}}\leq C_{q},$ (5.6)
and
$\|\alpha^{t,x,\phi^{r,\varepsilon}}\|_{\mathcal{H}^{q}},\left\|\frac{\delta\alpha^{t,x}}{\delta\phi_{i}}(\phi^{r,\varepsilon};\phi^{\prime}_{i})\right\|_{\mathcal{H}^{q}},\left\|\frac{\delta\alpha^{t,x}}{\delta\phi_{j}}(\phi^{r,\varepsilon};\phi^{\prime\prime}_{j})\right\|_{\mathcal{H}^{q}},\left\|\frac{\delta^{2}\alpha^{t,x}}{\delta\phi_{i}\phi_{j}}(\phi^{r,\varepsilon};\phi^{\prime}_{i},\phi^{\prime\prime}_{j})\right\|_{\mathcal{H}^{q}}\leq C_{q}.$ (5.7)
Hence, the uniform moment estimate (5.6) follows by applying standard moment estimates of SDEs (see [47, Theorem 3.4.3]) to (4.5), (4.6) and (4.7). The estimate (5.7) then follows from (4.9), (4.10) and the Cauchy-Schwarz inequality. This verifies Condition 1.
from which one can deduce uniform moment estimates of the state and control processes as in (5.6) and (5.7). As $\lim_{\varepsilon_{i}\downarrow 0}\phi^{\varepsilon}_{i}=\phi^{0}_{i}$ a.e., by stability results of SDEs (e.g., [47, Theorem 3.4.2]),
5.4 Proof of Theorem 4.5
@@
5.5 Proofs of Theorem 4.6, Corollary 4.7 and Proposition 4.8
Proof of Theorem 4.6.
which along with (4.32b) and (4.33b) shows that
where we used the symmetry of $\Psi_{i}(\cdot)$. Then by (4.28),
$\begin{split}&\frac{\delta V^{t,x}_{i}}{\delta\phi_{h}}(\phi;\phi^{\prime}_{h})-{\frac{\delta\overline{V}^{\phi}_{i}}{\delta\phi_{h}}}(t,x;\phi^{\prime}_{h})\\ &\quad={\mathbb{E}}\bigg{[}\int_{t}^{T}\bigg{(}X_{s}^{\top}Q_{i}Y^{h}_{s}+(KX_{s})^{\top}R_{i}KY^{h}_{s}-X_{s}^{\top}\Psi_{i}B_{h}K^{\prime}_{h}X_{s}\bigg{)}{\mathrm{d}}s+X_{T}^{\top}G_{i}Y^{h}_{T}\bigg{]}.\end{split}$ (5.18)
This along with (5.18) proves $\frac{\delta V^{t,x}_{i}}{\delta\phi_{h}}(\phi;\phi^{\prime}_{h})-{\frac{\delta\overline{V}^{\phi}_{i}}{\delta\phi_{h}}}(t,x;\phi^{\prime}_{h})=0$.
$\begin{split}\frac{\delta^{2}V^{t,x}_{i}}{\delta\phi_{h}\delta\phi_{\ell}}(\phi;\phi^{\prime}_{h},\phi^{\prime\prime}_{\ell})&={\mathbb{E}}\bigg{[}\int_{t}^{T}\bigg{(}Y^{\ell}_{s}Q_{i}Y^{h}_{s}+(KY^{\ell}_{s})^{\top}R_{i}KY^{h}_{s}\\ &\quad+(KY^{h}_{s})^{\top}(R_{i})_{\ell}K^{\prime\prime}_{\ell}X_{s}+(KY^{\ell}_{s})^{\top}(R_{i})_{h}K^{\prime}_{h}X_{s}+(K^{\prime\prime}_{\ell}X_{s})^{\top}(R_{i})_{\ell h}K^{\prime}_{h}X_{s}\\ &\quad+X^{\top}_{s}Q_{i}Z_{s}+(KX_{s})^{\top}\Big{(}R_{i}KZ_{s}+(R_{i})_{\ell}K^{\prime\prime}_{\ell}Y^{h}_{s}+(R_{i})_{h}K^{\prime}_{h}Y^{\ell}_{s}\Big{)}\bigg{)}{\mathrm{d}}s\bigg{]}\\ &\quad+{\mathbb{E}}\left[(Y^{\ell}_{T})^{\top}G_{i}Y^{h}_{T}+X^{\top}_{T}G_{i}Z_{T}\right].\end{split}$ (5.19)
${\mathbb{E}}\left[(Y^{\ell}_{T})^{\top}G_{i}Y^{h}_{T}+X^{\top}_{T}G_{i}Z_{T}\right]$
Substituting this into (5.19) implies that
$\begin{split}&\frac{\delta^{2}V^{t,x}_{i}}{\delta\phi_{h}\delta\phi_{\ell}}(\phi;\phi^{\prime}_{h},\phi^{\prime\prime}_{\ell})\\ &\quad={\mathbb{E}}\bigg{[}\int_{t}^{T}\bigg{(}(Y^{\ell}_{s})^{\top}\left(\Psi_{i}B_{h}K^{\prime}_{h}+(B_{h}K^{\prime}_{h})^{\top}\Psi_{i}+K^{\top}(R_{i})_{h}K^{\prime}_{h}+(K^{\prime}_{h})^{\top}((R_{i})_{h})^{\top}K\right)X_{s}\\ &\qquad+(Y^{h}_{s})^{\top}\left(\Psi_{i}B_{\ell}K^{\prime\prime}_{\ell}+(B_{\ell}K^{\prime\prime}_{\ell})^{\top}\Psi_{i}+K^{\top}(R_{i})_{\ell}K^{\prime\prime}_{\ell}+(K^{\prime\prime}_{\ell})^{\top}((R_{i})_{\ell})^{\top}K\right)X_{s}\\ &\qquad+(K^{\prime\prime}_{\ell}X_{s})^{\top}(R_{i})_{\ell h}K^{\prime}_{h}X_{s}\bigg{)}{\mathrm{d}}s\bigg{]}.\end{split}$ (5.20)
$\begin{split}{\frac{\delta^{2}\overline{V}^{\phi}_{i}}{\delta\phi_{h}\delta\phi_{\ell}}}(t,x;\phi^{\prime}_{h},\phi^{\prime\prime}_{\ell})&={\mathbb{E}}\left[\int_{t}^{T}\left(X_{s}^{\top}\Theta^{h}_{i}B_{\ell}K^{\prime\prime}_{\ell}X_{s}+X_{s}^{\top}\Theta^{\ell}_{i}B_{h}K^{\prime}_{h}X_{s}+(K^{\prime\prime}_{\ell}X_{s})^{\top}(R_{i})_{\ell h}K^{\prime}_{h}X_{s}\right){\mathrm{d}}s\right].\end{split}$ (5.21)
which along with (4.33b), (5.20) and (5.21) implies that $\frac{\delta^{2}V^{t,x}_{i}}{\delta\phi_{h}\delta\phi_{\ell}}(\phi;\phi^{\prime}_{h},\phi^{\prime\prime}_{\ell})={\frac{\delta^{2}\overline{V}^{\phi}_{i}}{\delta\phi_{h}\delta\phi_{\ell}}}(t,x;\phi^{\prime}_{h},\phi^{\prime\prime}_{\ell})$.
@@
Appendix A Proofs of Technical Lemmas
A.1 Proofs of Lemmas 4.1 and 4.2

### v3

@@ 1 Introduction
Throughout the paper, unless otherwise stated, proofs of theorems are deferred to Section 5.
@@ 3 Distributed Game as Markov Potential Game
This crowd motion game has been studied in [4] for pedestrian motion modelling, in [10] for animal flocking, and [9] for synchronization in neuroscience. We emphasize that the form of game here generalizes those in [4, 10, 9]: condition (3.5) removes the homogeneous and the symmetric interaction conditions imposed in their analysis. In particular, it allows each player to have different individual preferences specified by $c_{i}$ and $\ell_{i}$, as well as to be differently influenced by other players’ positions as indicated by the kernels $h_{ij}$ and $r_{ij}$.
@@ 4 Characterization of Differentiable Potential Game
We then present two lemmas regarding the linear derivative of $f:\pi^{(N)}\rightarrow{\mathbb{R}}$, which are crucial for Theorem 4.3. Lemma 4.1 shows that if $f$ has a linear derivative in $\pi_{i}$, then $f$ is differentiable along any line segment within $\pi_{i}$. Additionally, Lemma 4.2 shows that if $f$ is differentiable concerning all unilateral deviations, then $f$ is differentiable with respect to perturbations of all players’ policies, and the derivative can be computed via a chain rule. The proofs are given in Appendix B.2.
@@
We now show that if the value functions of a dynamics game are sufficiently regular and have a symmetric Jacobian, then this game is a potential game, and its dynamic potential function can be constructed via the linear derivative of its value function.
@@ 4.1 Probabilistic Characterization of Continuous-time Potential Game
where $B^{\phi}(t,x)\coloneqq B(t,x,\phi(t,x))$, $(\partial_{a_{i}}B^{\phi})[\phi^{\prime}_{i}](t,x)\coloneqq(\partial_{a_{i}}B)(t,x,\phi(t,x))\phi^{\prime}_{i}(t,x)$, and $\Sigma^{\phi}(t,x)$ and $(\partial_{a_{i}}\Sigma^{\phi})[\phi^{\prime}_{i}](t,x)$ are defined similarly. Here the differentiations are taken componentwise, i.e., $((\partial_{x}B^{\phi})(s,X^{t,x,\phi}_{s})Y_{s})_{\ell}=\sum_{j=1}^{n_{x}}\partial_{x_{j}}B^{\phi}_{\ell}(s,X^{t,x,\phi}_{s})(Y_{s})_{j}$ for all $\ell=1,\ldots,n_{x}$. Equation (4.6) is the sensitivity equation of (4.5) with respect to $\phi_{i}$ (see Lemma A.1).
@@
and $F_{\Sigma}(\cdot,y_{1},y_{2},\phi^{\prime}_{i},\phi^{\prime\prime}_{j})$ is defined similarly. Equation (4.7) is the sensitivity equation of (4.5) with respect to $\phi_{i}$ and $\phi_{j}$ (see Lemma A.2).
@@
To prove Theorem 4.4, one first shows that (4.6) and (4.7) characterize the derivatives of $X^{t,x,\phi}$ with respect to policies, using the pathwise differentiability of solutions to SDEs [25, Theorem 4, p. 105]. Then, by the regularity assumptions and the chain rule [25, Theorem 9, p. 97], one can show that (4.9) and (4.10) characterize the derivatives of $\alpha^{t,x,\phi}$ with respect to policies, and further establish (4.11) and (2). The detailed proof is given in Appendix A.1.
@@ 4.2 PDE Characterization of Continuous-time Potential Games
4.3 Continuous-time Linear Quadratic Potential Game
@@
The following theorem proves the equivalence of Theorems 4.4 and 4.5 for LQ games. The proof follows from a direct application of Itô’s formula, whose detail is given in Appendix A.2.
@@ 5 Proofs of Main Results
5.4 Proofs of Corollary 4.7 and Proposition 4.8
@@
Appendix A Proofs of Theorems 4.4 and 4.6
A.1 Proof of Theorem 4.4
We start by presenting Lemmas A.1, A.2 and A.3, which characterize the derivatives of $X^{t,x,\phi}$ and $\alpha^{t,x,\phi}$ with respect to policies. The proofs are given in Appendix B.2.
Lemma A.1.
  1. (1)
  2. (2)
Lemma A.2.
  1. (1)
  2. (2)
  3. (3)
Lemma A.3.
  1. (1)
  2. (2)
Proofs of Theorem 4.4 Items 1 and 2.
Fix $\phi\in\pi^{(N)}$. The linearity of $\frac{\delta V^{t,x}_{i}}{\delta\phi_{j}}(\phi;\cdot):\operatorname*{span}(\pi_{j})\rightarrow{\mathbb{R}}$ follows from the linearity of $\frac{\delta X^{t,x}}{\delta\phi_{j}}(\phi;\cdot)$ and $\frac{\delta\alpha^{t,x}}{\delta\phi_{j}}(\phi;\cdot)$ shown in Lemmas A.1 and A.3. Let $\phi^{\prime}_{j}\in\pi_{j}$ and for each $\varepsilon\in[0,1)$, let $\phi^{\varepsilon}=(\phi_{j}+\varepsilon(\phi^{\prime}_{j}-\phi_{j}),\phi_{-j})$. Recall that
Then [25, Theorem 9, p. 97], the $\mathcal{S}^{\infty}$-differentiability of $X^{t,x,\phi^{\varepsilon}}$ and the $\mathcal{H}^{\infty}$-differentiability of $\alpha^{t,x,\phi^{\varepsilon}}$ (see Lemmas A.1, A.2 and A.3) imply that $\frac{{\mathrm{d}}}{{\mathrm{d}}\varepsilon}V^{t,x}_{i}(\phi^{\varepsilon})|_{\varepsilon=0}=\frac{\delta V^{t,x}_{i}}{\delta\phi_{j}}(\phi;\phi^{\prime}_{j}-\phi_{j})$, hence Theorem 4.4 Item 1.
Proof of Theorem 4.4 Item 3.
$\|X^{t,x,\phi^{r,\varepsilon}}\|_{\mathcal{S}^{q}},\left\|\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi^{r,\varepsilon};\phi^{\prime}_{i})\right\|_{\mathcal{S}^{q}},\left\|\frac{\delta X^{t,x}}{\delta\phi_{j}}(\phi^{r,\varepsilon};\phi^{\prime\prime}_{j})\right\|_{\mathcal{S}^{q}},\left\|\frac{\delta^{2}X^{t,x}}{\delta\phi_{i}\phi_{j}}(\phi^{r,\varepsilon};\phi^{\prime}_{i},\phi^{\prime\prime}_{j})\right\|_{\mathcal{S}^{q}}\leq C_{q},$ (A.1)
and
$\|\alpha^{t,x,\phi^{r,\varepsilon}}\|_{\mathcal{H}^{q}},\left\|\frac{\delta\alpha^{t,x}}{\delta\phi_{i}}(\phi^{r,\varepsilon};\phi^{\prime}_{i})\right\|_{\mathcal{H}^{q}},\left\|\frac{\delta\alpha^{t,x}}{\delta\phi_{j}}(\phi^{r,\varepsilon};\phi^{\prime\prime}_{j})\right\|_{\mathcal{H}^{q}},\left\|\frac{\delta^{2}\alpha^{t,x}}{\delta\phi_{i}\phi_{j}}(\phi^{r,\varepsilon};\phi^{\prime}_{i},\phi^{\prime\prime}_{j})\right\|_{\mathcal{H}^{q}}\leq C_{q}.$ (A.2)
Hence, the uniform moment estimate (A.1) follows by applying standard moment estimates of SDEs (see [47, Theorem 3.4.3]) to (4.5), (4.6) and (4.7). The estimate (A.2) then follows from (4.9), (4.10) and the Cauchy-Schwarz inequality. This verifies Condition 1.
from which one can deduce uniform moment estimates of the state and control processes as in (A.1) and (A.2). As $\lim_{\varepsilon_{i}\downarrow 0}\phi^{\varepsilon}_{i}=\phi^{0}_{i}$ a.e., by stability results of SDEs (e.g., [47, Theorem 3.4.2]),
A.2 Proof of Theorem 4.6
Proof of Theorem 4.6.
which along with (4.32b) and (4.33b) shows that
where we used the symmetry of $\Psi_{i}(\cdot)$. Then by (4.28),
$\begin{split}&\frac{\delta V^{t,x}_{i}}{\delta\phi_{h}}(\phi;\phi^{\prime}_{h})-{\frac{\delta\overline{V}^{\phi}_{i}}{\delta\phi_{h}}}(t,x;\phi^{\prime}_{h})\\ &\quad={\mathbb{E}}\bigg{[}\int_{t}^{T}\bigg{(}X_{s}^{\top}Q_{i}Y^{h}_{s}+(KX_{s})^{\top}R_{i}KY^{h}_{s}-X_{s}^{\top}\Psi_{i}B_{h}K^{\prime}_{h}X_{s}\bigg{)}{\mathrm{d}}s+X_{T}^{\top}G_{i}Y^{h}_{T}\bigg{]}.\end{split}$ (A.3)
This along with (A.3) proves $\frac{\delta V^{t,x}_{i}}{\delta\phi_{h}}(\phi;\phi^{\prime}_{h})-{\frac{\delta\overline{V}^{\phi}_{i}}{\delta\phi_{h}}}(t,x;\phi^{\prime}_{h})=0$.
$\begin{split}\frac{\delta^{2}V^{t,x}_{i}}{\delta\phi_{h}\delta\phi_{\ell}}(\phi;\phi^{\prime}_{h},\phi^{\prime\prime}_{\ell})&={\mathbb{E}}\bigg{[}\int_{t}^{T}\bigg{(}Y^{\ell}_{s}Q_{i}Y^{h}_{s}+(KY^{\ell}_{s})^{\top}R_{i}KY^{h}_{s}\\ &\quad+(KY^{h}_{s})^{\top}(R_{i})_{\ell}K^{\prime\prime}_{\ell}X_{s}+(KY^{\ell}_{s})^{\top}(R_{i})_{h}K^{\prime}_{h}X_{s}+(K^{\prime\prime}_{\ell}X_{s})^{\top}(R_{i})_{\ell h}K^{\prime}_{h}X_{s}\\ &\quad+X^{\top}_{s}Q_{i}Z_{s}+(KX_{s})^{\top}\Big{(}R_{i}KZ_{s}+(R_{i})_{\ell}K^{\prime\prime}_{\ell}Y^{h}_{s}+(R_{i})_{h}K^{\prime}_{h}Y^{\ell}_{s}\Big{)}\bigg{)}{\mathrm{d}}s\bigg{]}\\ &\quad+{\mathbb{E}}\left[(Y^{\ell}_{T})^{\top}G_{i}Y^{h}_{T}+X^{\top}_{T}G_{i}Z_{T}\right].\end{split}$ (A.4)
${\mathbb{E}}\left[(Y^{\ell}_{T})^{\top}G_{i}Y^{h}_{T}+X^{\top}_{T}G_{i}Z_{T}\right]$
Substituting this into (A.4) implies that
$\begin{split}&\frac{\delta^{2}V^{t,x}_{i}}{\delta\phi_{h}\delta\phi_{\ell}}(\phi;\phi^{\prime}_{h},\phi^{\prime\prime}_{\ell})\\ &\quad={\mathbb{E}}\bigg{[}\int_{t}^{T}\bigg{(}(Y^{\ell}_{s})^{\top}\left(\Psi_{i}B_{h}K^{\prime}_{h}+(B_{h}K^{\prime}_{h})^{\top}\Psi_{i}+K^{\top}(R_{i})_{h}K^{\prime}_{h}+(K^{\prime}_{h})^{\top}((R_{i})_{h})^{\top}K\right)X_{s}\\ &\qquad+(Y^{h}_{s})^{\top}\left(\Psi_{i}B_{\ell}K^{\prime\prime}_{\ell}+(B_{\ell}K^{\prime\prime}_{\ell})^{\top}\Psi_{i}+K^{\top}(R_{i})_{\ell}K^{\prime\prime}_{\ell}+(K^{\prime\prime}_{\ell})^{\top}((R_{i})_{\ell})^{\top}K\right)X_{s}\\ &\qquad+(K^{\prime\prime}_{\ell}X_{s})^{\top}(R_{i})_{\ell h}K^{\prime}_{h}X_{s}\bigg{)}{\mathrm{d}}s\bigg{]}.\end{split}$ (A.5)
$\begin{split}{\frac{\delta^{2}\overline{V}^{\phi}_{i}}{\delta\phi_{h}\delta\phi_{\ell}}}(t,x;\phi^{\prime}_{h},\phi^{\prime\prime}_{\ell})&={\mathbb{E}}\left[\int_{t}^{T}\left(X_{s}^{\top}\Theta^{h}_{i}B_{\ell}K^{\prime\prime}_{\ell}X_{s}+X_{s}^{\top}\Theta^{\ell}_{i}B_{h}K^{\prime}_{h}X_{s}+(K^{\prime\prime}_{\ell}X_{s})^{\top}(R_{i})_{\ell h}K^{\prime}_{h}X_{s}\right){\mathrm{d}}s\right].\end{split}$ (A.6)
which along with (4.33b), (A.5) and (A.6) implies that $\frac{\delta^{2}V^{t,x}_{i}}{\delta\phi_{h}\delta\phi_{\ell}}(\phi;\phi^{\prime}_{h},\phi^{\prime\prime}_{\ell})={\frac{\delta^{2}\overline{V}^{\phi}_{i}}{\delta\phi_{h}\delta\phi_{\ell}}}(t,x;\phi^{\prime}_{h},\phi^{\prime\prime}_{\ell})$.
Appendix B Proofs of Technical Lemmas
B.1 Proofs of Lemmas 4.1 and 4.2